# Forget BIT, It is All about TOKEN: Towards Semantic Information Theory for LLMs

### v1

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities in numerous real-world applications. While the vast majority of research conducted from an experimental perspective is progressing rapidly, it demands substantial computational power, data, and other resources. Therefore, how to open the black-box of LLMs from a theoretical standpoint has become a critical challenge. This paper takes the theory of rate-distortion function, directed information, and Granger causality as its starting point to investigate the information-theoretic principles behind LLMs, leading to the development of semantic information theory for LLMs, where the fundamental unit is token, rather than bits that lacks any semantic meaning. By defining the probabilistic model of LLMs, we discuss structure-agnostic information-theoretic measures, such as the directed rate-distortion function in pre-training, the directed rate-reward function in post-training, and the semantic information flow in inference phase. This paper also delves deeply into the theory of token-level semantic embedding and the information-theoretically optimal vectorization method. Thereafter, we propose a general definition of autoregression LLM, where the Transformer architecture and its performance such as ELBO, generalization error bound, memory capacity, and semantic information measures can be derived theoretically. Other architectures, such as Mamba/Mamba2 and LLaDA, are also discussed in our framework. Consequently, this paper provides a theoretical framework for understanding LLMs from the perspective of semantic information theory, which also offers the necessary theoretical tools for further in-depth research.
@@ II-C Granger Causality
Let $\mathcal{U}_{t}$ be all the knowledge in the universe available at time $t$ with $1\leq t\leq n$, $\mathcal{U}_{t}^{-}$ be the knowledge in the modified universe in which $X_{1:n}$ is excluded, $X_{t}$ is said to cause $Y_{t+1}$ if
$P(Y_{t+1}\in\mathcal{A}|\mathcal{U}_{t})\neq P(Y_{t+1}\in\mathcal{A}|\mathcal{U}_{t}^{-}).$ (6)
@@ III LLM as a Next-token Predictor
Inspired from information theory, this section will introduce the probabilistic model and architecture irrelevant properties for LLMs.
@@ III-A Probabilistic Model of LLMs
Figure 1: The probabilistic model of an LLM at time $t\in\mathbb{N}$, where $X_{1:n}$ is the token sequence with $1\leq n<t\leq T$ whose semantic vector embedding is $S_{1:n}$, $Y_{n+1:T}$ is the output token sequence, whose semantic vector embedding is $U_{n+1:T}$. $\Phi$ represents the parameters after training.
The probabilistic model of LLMs is illustrated in Fig. 1. The input token sequence is $X_{1:n}$ with $1\leq n<t\leq T$ and $n\in\mathbb{N}$, which will be mapped to semantic vector sequence $S_{1:n}$ by a semantic embedding module $f$. The LLM is modeled as a transition probability with a parameter $\Phi$, which represents the parameters of the LLM after training. The LLM generates the embedding of next token $U_{t}$ based on $S_{1:n}$ and the previously generated $U_{n+1:t-1}$, that is
$P(U_{t}|U_{n+1:t-1},S_{1:n};\Phi).$ (7)
$\varphi$ is an inverse module of embedding, which maps $U_{t}$ to the output token $Y_{t}$. It should be noticed that the probabilistic model in Fig. 1 is general and architecture irrelevant.
@@
$K(\mathbf{y})=\min_{i,\mathbf{p}}\{K(i)+l(\mathbf{p}):T_{i}(\mathbf{p})=\mathbf{y}\}$ (9)
where $i\in\mathbb{N}$ is the index of a sequence of Turing machines. It can be seen that $K(\mathbf{y})$ is decomposed into two parts: the first part is a Turing machine $T_{i}$, i.e., the meaningful information or model in the data, and the second part is the irregular aspects of $\mathbf{y}$, i.e., a program $\mathbf{p}$ to be interpreted by $T_{i}$. Following this idea, the LLM is equivalent to $T_{i}$, $\mathbf{p}$ is the input $\mathbf{x}$, i.e., the prompt.
@@ III-B Directed Rate-distortion Function in Pre-training Phase
From the perspective of information theory, Eq. (7) is a discrete-time channel with feedback and state [42, 64]. The input is $S_{1:n}$, the output is $U_{n+1:T}$, the feedback at time $t$ is $U_{t}$, and the channel state is the parameter $\Phi$. In contrast to the reliable communication problem, the goal here is to ensure the output token sequence aligns with our expectations, rather than a flawless recovery of the input.
As a discrete-time channel with feedback and state, the directed information is a natural choice to measure the information transferred from $S_{1:n}$ to $U_{n+1:T}$ with parameter $\Phi$ [42, 64]. According to Definition 2, we have
$I(S_{1:n}\to U_{n+1:T};\Phi)=\sum_{t=n+1}^{T}I(S_{1:n};U_{t}|U_{n+1:t-1};\Phi).$ (10)
Let $U_{n+1:T}^{\hbar}$ be the labeled sequence by human being with the input $S_{1:n}$, and $D_{KL}(\cdot\|\cdot)$ be the KL divergence. Denote $P_{t}^{\hbar}=P(U_{t}^{\hbar}|U_{n+1:t-1}^{\hbar},S_{1:n})$ and $Q_{t}^{\Phi}=P(U_{t}|U_{n+1:t-1},S_{1:n};\Phi)$ for $t=n+1,\ldots,T$, we then have the following definition.
@@
$R_{pre}(D)=\frac{1}{T}\inf_{\Phi:\frac{1}{T}\sum_{t=n+1}^{T}D_{KL}(P_{t}^{\hbar}\|Q_{t}^{\Phi})<D}I(S_{1:n}\to U_{n+1:T};\Phi).$ (11)
Similar to Shannon capacity, $R_{pre}(D)$ is defined as a universal measure connecting the input sequence $S_{1:n}$ and output sequence $U_{n+1:T}$. Furthermore, $R_{pre}(D)$ is independent of any implementation methods, such as Transformer or novel architectures yet to be conceived. In contrast to the classical rate-distortion function in Definition 1, which governs a lossy source codecs, the output sequence $U_{n+1:T}$ in this context is instead constrained by a condition defined in terms of KL divergence. Therefore, $R_{pre}(D)$ is the minimum information needed from $S_{1:n}$ to generate the expected $U_{n+1:T}$ with an average distortion $D$. The curve of $R_{pre}(D)$ versus the optimization process of $\Phi$ will reveal key properties of the pre-training in practice. Simple derivation will give us the following theorem.
@@
The cross-entropy between $P_{t}^{\hbar}$ and $Q_{t}^{\Phi}$ is given by
$H(P_{t}^{\hbar},Q_{t}^{\Phi})=H(P_{t}^{\hbar})+D_{KL}(P_{t}^{\hbar}\|Q_{t}^{\Phi}),\quad t=n+1,\ldots,T.$ (13)
@@
$\min_{\Phi}H(P_{t}^{\hbar},Q_{t}^{\Phi})\Leftrightarrow\min_{\Phi}D_{KL}(P_{t}^{\hbar}\|Q_{t}^{\Phi}),\quad t=n+1,\ldots,T.$ (14)
@@
$Q_{t}^{\Phi^{\hbar}}=P(U_{t}|U_{n+1:t-1},S_{1:n};\Phi^{\hbar})=P(U_{t}^{\hbar}|U_{n+1:t-1}^{\hbar},S_{1:n})=P_{t}^{\hbar},\quad t=n+1,\ldots,T,$ (15)
@@
$I(S_{1:n}\to U_{n+1:T};\Phi^{\hbar})=\sum_{t=n+1}^{T}I(S_{1:n};U_{t}|U_{n+1:t-1};\Phi^{\hbar})$ (17)
$=$ $\sum_{t=n+1}^{T}H(U_{t}|U_{n+1:t-1};\Phi^{\hbar})-\sum_{t=n+1}^{T}H(U_{t}|U_{n+1:t-1},S_{1:n};\Phi^{\hbar})$
$=$ $\sum_{t=n+1}^{T}H(U_{t}^{\hbar}|U_{n+1:t-1}^{\hbar})-\sum_{t=n+1}^{T}H(U_{t}^{\hbar}|U_{n+1:t-1}^{\hbar},S_{1:n})$
$=$ $I(S_{1:n}\to U_{n+1:T}^{\hbar}).$
@@
$\mathcal{L}(\Phi)=I(S_{1:n};U_{t}|U_{n+1:t-1};\Phi)+\lambda H(P_{t}^{\hbar},Q_{t}^{\Phi}),\quad t=n+1,\ldots,T,$ (18)
@@
Consider the pre-training phase, the distribution before and after one training step is denoted by $P_{t}(\Phi)=P(U_{t}|U_{n+1:t-1},S_{1:n};\Phi)$ and $P_{t}(\Phi^{\prime})=P(U_{t}|U_{n+1:t-1},S_{1:n};\Phi^{\prime})$, respectively. According to [76], the entry of the Fisher information matrix at the $i$-th row and $j$-th column is given by
$\mathcal{I}(\Phi)]_{ij}=\left.\frac{\partial^{2}}{\partial\Phi^{\prime}_{i}\partial\Phi^{\prime}_{j}}H(P_{t}(\Phi),P_{t}(\Phi^{\prime}))\right|_{\Phi^{\prime}=\Phi}.$ (19)
Thus, the Fisher information matrix represents the curvature of the cross-entropy loss with respect to the parameters $\Phi$. By modifying the gradient with Fisher information matrix, the natural gradient method is then proposed for neural network training. Due to the high computation complexity and storage cost, the Kronecker-factored approximate curvature method is used in practice [77].
@@ III-C Directed Rate-reward Function in Post-training Phase
An evaluation function $w(S_{1:n},U_{n+1:T})$, the reward function in RLHF, is introduced to assign a score to the generated sequence $U_{n+1:T}$ for the input sequence $S_{1:n}$. We then have the following definition.
@@
$R_{post}(W)=\frac{1}{T}\inf_{\Phi^{\hbar}:w(S_{1:n},U_{n+1:T})>W}I(S_{1:n}\to U_{n+1:T};\Phi^{\hbar}).$ (20)
@@
$\mathcal{L}(\Phi^{\hbar})=I(S_{1:n}\to U_{n+1:T};\Phi^{\hbar})-\lambda w(S_{1:n},U_{n+1:T}),$ (21)
@@
Theorem 1 shows that the LLM approaches $I(S_{1:n}\to U_{n+1:T}^{\hbar})$ during pre-training, which measures the information transferred from $S_{1:n}$ to $U_{n+1:T}^{\hbar}$ by human being. The post-training further adjusts the parameter from $\Phi^{\hbar}$ to $\Phi^{\hbar+}$ such that the generated sequence $U_{n+1:T}$ meets human preferences. Recalling the discussion in Section II-C, we have the following conclusion.
@@ III-D Semantic Information Flow in Inference Phase
During the inference phase, the LLM with parameter $\Phi^{\hbar+}$ is employed to generate the output token sequence $U_{n+1:T}$ based on the input token sequence $S_{1:n}$. In contrast to the post-training phase, where the focus is on the average performance across all possible output sequences, the inference phase considers the specific output sequence for the given input sequence. Therefore, it is natural to use the directed information density in Definition 4 to analyze the inference process. The semantic information flow can then be defined as follows.
@@
The semantic information flow for LLMs is defined as the directed information density from $S_{1:n}$ to $U_{n+1:t}$ as follows:
$\imath(S_{1:n}\to U_{n+1:t};\Phi^{\hbar+})=\sum_{\tau=n+1}^{t}\imath(S_{1:n};U_{\tau}|U_{n+1:\tau-1};\Phi^{\hbar+}),\quad t=n+1,\ldots,T.$ (22)
@@
The semantic information flow $\imath(S_{1:n}\to U_{n+1:t};\Phi^{\hbar+})$ is a Markovian sub-martingale for $t=n+1,\ldots,T$.
@@
$\imath(S_{1:n}\to U_{n+1:t};\Phi^{\hbar+})=\imath(S_{1:n}\to U_{n+1:t-1};\Phi^{\hbar+})+\imath(S_{1:n};U_{t}|U_{n+1:t-1};\Phi^{\hbar+}),$ (23)
@@
$\imath(S_{1:n};U_{t}|U_{n+1:t-1};\Phi^{\hbar+})=\log\frac{P(U_{t}|U_{n+1:t-1},S_{1:n};\Phi^{\hbar+})}{P(U_{t}|U_{n+1:t-1};\Phi^{\hbar+})}.$ (24)
@@
$\mathbb{E}\{\imath(S_{1:n}\to U_{n+1:t};\Phi^{\hbar+})|\imath(S_{1:n}\to U_{n+1:t-1};\Phi^{\hbar+}),\ldots,\imath(S_{1:n}\to U_{n+1};\Phi^{\hbar+})\}$ (25)
$=$ $\mathbb{E}\{\imath(S_{1:n}\to U_{n+1:t};\Phi^{\hbar+})|\imath(S_{1:n}\to U_{n+1:t-1};\Phi^{\hbar+})\}$
$=$ $\imath(S_{1:n}\to U_{n+1:t-1};\Phi^{\hbar+})+\mathbb{E}\{\imath(S_{1:n};U_{t}|U_{n+1:t-1};\Phi^{\hbar+})\}$
$=$ $\imath(S_{1:n}\to U_{n+1:t-1};\Phi^{\hbar+})+D_{KL}(P(U_{t}|U_{n+1:t-1},S_{1:n};\Phi^{\hbar+})\|P(U_{t}|U_{n+1:t-1};\Phi^{\hbar+}))$
$\geq$ $\imath(S_{1:n}\to U_{n+1:t-1};\Phi^{\hbar+}).$
@@
In the following, we will discuss the properties of semantic information flow as a sub-martingale. According to Doob decomposition, we have
$\imath(S_{1:n}\to U_{n+1:t};\Phi^{\hbar+})=M_{t}+A_{t},$ (26)
where $A_{t}$ is a predictable and non-decreasing process
$A_{t}=\sum_{j=n+1}^{t}\mathbb{E}\{\imath(S_{1:n}\to U_{n+1:j};\Phi^{\hbar+})-\imath(S_{1:n}\to U_{n+1:j-1};\Phi^{\hbar+})|\imath(S_{1:n}\to U_{n+1:j-1};\Phi^{\hbar+})\}$ (27)
and $M_{t}$ is a martingale
$M_{t}=$ $\imath(S_{1:n}\to U_{n+1};\Phi^{\hbar+})$ (28)
$+\sum_{j=n+2}^{t}(\imath(S_{1:n}\to U_{n+1:j};\Phi^{\hbar+})-\imath(S_{1:n}\to U_{n+1:j-1};\Phi^{\hbar+})-A_{j}).$
@@
$\Pr\{M_{t}>\alpha,V_{t}<\beta\}\leq\exp\left(-\frac{\alpha^{2}}{2(\alpha+\beta)}\right).$ (30)

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities in numerous real-world applications. While the vast majority of research conducted from an experimental perspective is progressing rapidly, it demands substantial computational power, data, and other resources. Therefore, how to open the black-box of LLMs from a theoretical standpoint has become a critical challenge. This paper takes the theory of rate-distortion function, directed information, and Granger causality as its starting point to investigate the information-theoretic principles behind LLMs, leading to the development of semantic information theory for LLMs, where the fundamental unit is token, rather than bits that lacks any semantic meaning. By defining the probabilistic model of LLMs, we discuss architecture-agnostic information-theoretic measures, such as the directed rate-distortion function in pre-training, the directed rate-reward function in post-training, and the semantic information flow in inference phase. This paper also delves deeply into the theory of token-level semantic embedding and the information-theoretically optimal vectorization method. Thereafter, we propose a general definition of autoregression LLM, which can also be formulated as the Boltzmann distribution in statistical physics. The Transformer architecture and its performance such as ELBO, generalization error bound, memory capacity, and semantic information measures can then be derived theoretically. Other architectures, such as Mamba/Mamba2 and LLaDA, are also discussed in our framework. Consequently, this paper provides a theoretical framework for understanding LLMs from the perspective of semantic information theory, which also offers the necessary theoretical tools for further in-depth research.
@@ II-C Granger Causality
Let $\mathcal{U}_{i}$ be all the knowledge in the universe available at time $i$ with $1\leq i\leq n$, $\mathcal{U}_{i}^{-}$ be the knowledge in the modified universe in which $X_{1:n}$ is excluded, $X_{i}$ is said to cause $Y_{i+1}$ if
$P(Y_{i+1}\in\mathcal{A}|\mathcal{U}_{i})\neq P(Y_{i+1}\in\mathcal{A}|\mathcal{U}_{i}^{-}).$ (6)
@@ III LLM as a Next-token Predictor
Inspired from information theory, this section will introduce the probabilistic model and architecture-agnostic properties for LLMs.
@@ III-A Probabilistic Model of LLMs
Figure 1: The probabilistic model of an LLM at time $i\in\mathbb{N}$, where $X_{1:n}$ is the token sequence with $1\leq n<i\leq N$ whose semantic vector embedding is $S_{1:n}$, $Y_{n+1:N}$ is the output token sequence, whose semantic vector embedding is $U_{n+1:N}$. $\Phi$ represents the parameters after training.
The probabilistic model of LLMs is illustrated in Fig. 1. The input token sequence is $X_{1:n}$ with $1\leq n<i\leq N$ and $n\in\mathbb{N}$, which will be mapped to semantic vector sequence $S_{1:n}$ by a semantic embedding module $f$. The LLM is modeled as a transition probability with a parameter $\Phi$, which represents the parameters of the LLM after training. The LLM generates the embedding of next token $U_{i}$ based on $S_{1:n}$ and the previously generated $U_{n+1:i-1}$, that is
$P(U_{i}|U_{n+1:i-1},S_{1:n};\Phi).$ (7)
$\varphi$ is an inverse module of embedding, which maps $U_{i}$ to the output token $Y_{i}$. It should be noticed that the probabilistic model in Fig. 1 is general and architecture irrelevant.
@@
$K(\mathbf{y})=\min_{m,\mathbf{p}}\{K(m)+l(\mathbf{p}):T_{m}(\mathbf{p})=\mathbf{y}\}$ (9)
where $m\in\mathbb{N}$ is the index of a sequence of Turing machines. It can be seen that $K(\mathbf{y})$ is decomposed into two parts: the first part is a Turing machine $T_{m}$, i.e., the meaningful information or model in the data, and the second part is the irregular aspects of $\mathbf{y}$, i.e., a program $\mathbf{p}$ to be interpreted by $T_{m}$. Following this idea, the LLM is equivalent to $T_{m}$, $\mathbf{p}$ is the input $\mathbf{x}$, i.e., the prompt.
@@ III-B Directed Rate-distortion Function in Pre-training Phase
From the perspective of information theory, Eq. (7) is a discrete-time channel with feedback and state [42, 64]. The input is $S_{1:n}$, the output is $U_{n+1:N}$, the feedback at time $i$ is $U_{i}$, and the channel state is the parameter $\Phi$. In contrast to the reliable communication problem, the goal here is to ensure the output token sequence aligns with our expectations, rather than a flawless recovery of the input.
As a discrete-time channel with feedback and state, the directed information is a natural choice to measure the information transferred from $S_{1:n}$ to $U_{n+1:N}$ with parameter $\Phi$ [42, 64]. According to Definition 2, we have
$I(S_{1:n}\to U_{n+1:N};\Phi)=\sum_{i=n+1}^{N}I(S_{1:n};U_{i}|U_{n+1:i-1};\Phi).$ (10)
Let $U_{n+1:N}^{\hbar}$ be the labeled sequence by human being with the input $S_{1:n}$, and $D_{KL}(\cdot\|\cdot)$ be the KL divergence. Denote $P_{i}^{\hbar}=P(U_{i}^{\hbar}|U_{n+1:i-1}^{\hbar},S_{1:n})$ and $Q_{i}^{\Phi}=P(U_{i}|U_{n+1:i-1},S_{1:n};\Phi)$ for $i=n+1,\ldots,N$, we then have the following definition.
@@
$R_{pre}(D)=\frac{1}{N}\inf_{\Phi:\frac{1}{N}\sum_{i=n+1}^{N}D_{KL}(P_{i}^{\hbar}\|Q_{i}^{\Phi})<D}I(S_{1:n}\to U_{n+1:N};\Phi).$ (11)
Similar to Shannon capacity, $R_{pre}(D)$ is defined as a universal measure connecting the input sequence $S_{1:n}$ and output sequence $U_{n+1:N}$. Furthermore, $R_{pre}(D)$ is independent of any implementation methods, such as Transformer or novel architectures yet to be conceived. In contrast to the classical rate-distortion function in Definition 1, which governs a lossy source codecs, the output sequence $U_{n+1:N}$ in this context is instead constrained by a condition defined in terms of KL divergence. Therefore, $R_{pre}(D)$ is the minimum information needed from $S_{1:n}$ to generate the expected $U_{n+1:N}$ with an average distortion $D$. The curve of $R_{pre}(D)$ versus the optimization process of $\Phi$ will reveal key properties of the pre-training in practice. Simple derivation will give us the following theorem.
@@
The cross-entropy between $P_{i}^{\hbar}$ and $Q_{i}^{\Phi}$ is given by
$H(P_{i}^{\hbar},Q_{i}^{\Phi})=H(P_{i}^{\hbar})+D_{KL}(P_{i}^{\hbar}\|Q_{i}^{\Phi}),\quad i=n+1,\ldots,N.$ (13)
@@
$\min_{\Phi}H(P_{i}^{\hbar},Q_{i}^{\Phi})\Leftrightarrow\min_{\Phi}D_{KL}(P_{i}^{\hbar}\|Q_{i}^{\Phi}),\quad i=n+1,\ldots,N.$ (14)
@@
$Q_{i}^{\Phi^{\hbar}}=P(U_{i}|U_{n+1:i-1},S_{1:n};\Phi^{\hbar})=P(U_{i}^{\hbar}|U_{n+1:i-1}^{\hbar},S_{1:n})=P_{i}^{\hbar},\quad i=n+1,\ldots,N,$ (15)
@@
$I(S_{1:n}\to U_{n+1:N};\Phi^{\hbar})=\sum_{i=n+1}^{N}I(S_{1:n};U_{i}|U_{n+1:i-1};\Phi^{\hbar})$ (17)
$=$ $\sum_{i=n+1}^{N}H(U_{i}|U_{n+1:i-1};\Phi^{\hbar})-\sum_{i=n+1}^{N}H(U_{i}|U_{n+1:i-1},S_{1:n};\Phi^{\hbar})$
$=$ $\sum_{i=n+1}^{N}H(U_{i}^{\hbar}|U_{n+1:i-1}^{\hbar})-\sum_{i=n+1}^{N}H(U_{i}^{\hbar}|U_{n+1:i-1}^{\hbar},S_{1:n})$
$=$ $I(S_{1:n}\to U_{n+1:N}^{\hbar}).$
@@
$\mathcal{L}(\Phi)=I(S_{1:n};U_{i}|U_{n+1:i-1};\Phi)+\lambda H(P_{i}^{\hbar},Q_{i}^{\Phi}),\quad i=n+1,\ldots,N,$ (18)
@@
Consider the pre-training phase, the distribution before and after one training step is denoted by $P_{i}(\Phi)=P(U_{i}|U_{n+1:i-1},S_{1:n};\Phi)$ and $P_{i}(\Phi^{\prime})=P(U_{i}|U_{n+1:i-1},S_{1:n};\Phi^{\prime})$, respectively. According to [76], the entry of the Fisher information matrix at the $k$-th row and $l$-th column is given by
$\mathcal{I}(\Phi)]_{kl}=\left.\frac{\partial^{2}}{\partial\Phi^{\prime}_{k}\partial\Phi^{\prime}_{l}}H(P_{i}(\Phi),P_{i}(\Phi^{\prime}))\right|_{\Phi^{\prime}=\Phi}.$ (19)
Thus, the Fisher information matrix represents the curvature of the cross-entropy loss with respect to the parameters $\Phi$. By modifying the gradient with Fisher information matrix, the natural gradient method is then proposed for neural network training. Due to the high computation complexity and storage cost, the Kronecker-factored approximate curvature (K-FAC) method is used in practice [77].
@@ III-C Directed Rate-reward Function in Post-training Phase
An evaluation function $w(S_{1:n},U_{n+1:N})$, the reward function in RLHF, is introduced to assign a score to the generated sequence $U_{n+1:N}$ for the input sequence $S_{1:n}$. We then have the following definition.
@@
$R_{post}(W)=\frac{1}{N}\inf_{\Phi^{\hbar}:w(S_{1:n},U_{n+1:N})>W}I(S_{1:n}\to U_{n+1:N};\Phi^{\hbar}).$ (20)
@@
$\mathcal{L}(\Phi^{\hbar})=I(S_{1:n}\to U_{n+1:N};\Phi^{\hbar})-\lambda w(S_{1:n},U_{n+1:N}),$ (21)
@@
Theorem 1 shows that the LLM approaches $I(S_{1:n}\to U_{n+1:N}^{\hbar})$ during pre-training, which measures the information transferred from $S_{1:n}$ to $U_{n+1:N}^{\hbar}$ by human being. The post-training further adjusts the parameter from $\Phi^{\hbar}$ to $\Phi^{\hbar+}$ such that the generated sequence $U_{n+1:N}$ meets human preferences. Recalling the discussion in Section II-C, we have the following conclusion.
@@ III-D Semantic Information Flow in Inference Phase
During the inference phase, the LLM with parameter $\Phi^{\hbar+}$ is employed to generate the output token sequence $U_{n+1:N}$ based on the input token sequence $S_{1:n}$. In contrast to the training phase, where the focus is on the average performance across all possible output sequences, the inference phase considers the specific output sequence for the given input sequence. Therefore, it is natural to use the directed information density in Definition 4 to analyze the inference process. The semantic information flow can then be defined as follows.
@@
The semantic information flow for LLMs is defined as the directed information density from $S_{1:n}$ to $U_{n+1:i}$ as follows:
$\imath(S_{1:n}\to U_{n+1:i};\Phi^{\hbar+})=\sum_{j=n+1}^{i}\imath(S_{1:n};U_{j}|U_{n+1:j-1};\Phi^{\hbar+}),\quad i=n+1,\ldots,N.$ (22)
@@
The semantic information flow $\imath(S_{1:n}\to U_{n+1:i};\Phi^{\hbar+})$ is a Markovian sub-martingale for $i=n+1,\ldots,N$.
@@
$\imath(S_{1:n}\to U_{n+1:i};\Phi^{\hbar+})=\imath(S_{1:n}\to U_{n+1:i-1};\Phi^{\hbar+})+\imath(S_{1:n};U_{i}|U_{n+1:i-1};\Phi^{\hbar+}),$ (23)
@@
$\imath(S_{1:n};U_{i}|U_{n+1:i-1};\Phi^{\hbar+})=\log\frac{P(U_{i}|U_{n+1:i-1},S_{1:n};\Phi^{\hbar+})}{P(U_{i}|U_{n+1:i-1};\Phi^{\hbar+})}.$ (24)
@@
$\mathbb{E}\{\imath(S_{1:n}\to U_{n+1:i};\Phi^{\hbar+})|\imath(S_{1:n}\to U_{n+1:i-1};\Phi^{\hbar+}),\ldots,\imath(S_{1:n}\to U_{n+1};\Phi^{\hbar+})\}$ (25)
$=$ $\mathbb{E}\{\imath(S_{1:n}\to U_{n+1:i};\Phi^{\hbar+})|\imath(S_{1:n}\to U_{n+1:i-1};\Phi^{\hbar+})\}$
$=$ $\imath(S_{1:n}\to U_{n+1:i-1};\Phi^{\hbar+})+\mathbb{E}\{\imath(S_{1:n};U_{i}|U_{n+1:i-1};\Phi^{\hbar+})\}$
$=$ $\imath(S_{1:n}\to U_{n+1:i-1};\Phi^{\hbar+})+D_{KL}(P(U_{i}|U_{n+1:i-1},S_{1:n};\Phi^{\hbar+})\|P(U_{i}|U_{n+1:i-1};\Phi^{\hbar+}))$
$\geq$ $\imath(S_{1:n}\to U_{n+1:i-1};\Phi^{\hbar+}).$
@@
In the following, we will discuss the properties of semantic information flow as a sub-martingale. According to Doob’s decomposition, we have
$\imath(S_{1:n}\to U_{n+1:i};\Phi^{\hbar+})=M_{i}+A_{i},$ (26)
where $A_{i}$ is a predictable and non-decreasing process
$A_{i}=\sum_{j=n+1}^{i}\mathbb{E}\{\imath(S_{1:n}\to U_{n+1:j};\Phi^{\hbar+})-\imath(S_{1:n}\to U_{n+1:j-1};\Phi^{\hbar+})|\imath(S_{1:n}\to U_{n+1:j-1};\Phi^{\hbar+})\}$ (27)
and $M_{i}$ is a martingale
$M_{i}=$ $\imath(S_{1:n}\to U_{n+1};\Phi^{\hbar+})$ (28)
$+\sum_{j=n+2}^{i}(\imath(S_{1:n}\to U_{n+1:i};\Phi^{\hbar+})-\imath(S_{1:n}\to U_{n+1:i-1};\Phi^{\hbar+})-A_{j}).$
@@
$\Pr\{M_{i}>\alpha,V_{i}<\beta\}\leq\exp\left(-\frac{\alpha^{2}}{2(\alpha+\beta)}\right).$ (30)